# High-Throughput and Scalable Secure Inference Protocols for Deep Learning with Packed Secret Sharing

### v1

@@ VI-B Neural Networks & Datasets
    VI-C Communication Costs of Secure Inference
@@
    The communication costs of the protocol from Liu et al. [5] and our protocol on various datasets and neural networks are shown in Table I. Compared to Liu et al. [5], we achieve a significant advantage in terms of offline communication, online communication, and total communication costs. Moreover, our analysis reveals a progressively more pronounced advantage along the diagonal from the upper-left to the lower-right corner of the table. For instance, compared to Liu et al. [5], our approach improves the communication by upto $5.85\times$, $11.17\times$, and $6.83\times$ in offline, online, and total communication overhead respectively for AlexNet. Additionally, for VGG16, Liu et al. [5] suffers from memory overflow under both the 31PC where 12 parties are controlled by the adversary and 63PC, whereas our approach continues to execute successfully. This advantage primarily stems from the fact that more parties enable the selection of larger values for $k$ in PSS, where increased $k$ confers greater benefits for deeper neural networks and larger datasets.
    VI-D Running Time of Secure Inference
    VI-D1 WAN
@@
    The running time of Liu et al. [5] and ours in WAN is shown in Table II. Similar to Table I, our advantages become increasingly significant from the top-left to the bottom-right of the table. However, it is worth noting that in the case of 11PC, there is no improvement compared to Liu et al. [5]. This phenomenon primarily stems from our scheme is significantly affected by computational overhead, and when a small $k$ is selected, the communication efficiency improvement is limited. Furthermore, we can find that when the communication overhead increases by more than $3.80\times$, our online running time will be significantly improved, with a maximum increase of $2.61\times$. At the same time, in the case of more parties and a deeper neural network, the offline running time and the total running time can be increased by up to $1.52\times$ and $1.75\times$ respectively.
    VI-D2 LAN
@@
    The running time of Liu et al. [5] and ours in LAN is shown in Table III. We observe that in the online phase, our running time shows little to no advantage compared to Liu et al. [5], as our scheme is more heavily affected by computational overhead. However, the offline and the total running time, as the number of parties increases and the network deepens, will also have significant improvements, with the maximum increase being $3.76\times$ and $2.61\times$ respectively.
@@
    In a word, our improvements in the LAN setting are less pronounced compared to those in the WAN setting. The reason lies in the higher computational overhead required by our scheme. Fortunately, as the number of parties increases and the neural networks become deeper, our scheme exhibits more substantial advantages, mainly owing to our efficient vector-matrix multiplication, parallel convolution computation, and parallel non-linear computation based on PSS. Moreover, MPC protocols are more commonly deployed in WAN settings, so our improvements hold practical significance.
@@ VII Related Work
    Secure neural network inference has been a prominent research focus in recent years. Based on the number of participating parties, existing works can be broadly categorized into two categories: those involving a small number of parties (e.g., 2PC, 3PC, 4PC) and those designed for a large number of parties (nnPC).
    In the area of two-party computation (2PC), ABY [20] and its follow-up work ABY2.0 [21] design a framework that combines secure computation schemes based on arithmetic sharing, boolean sharing, and Yao’s garbled circuits (GC). SecureML [17] proposes a scheme in the two-server outsourced computation models with secret sharing and GC. Meanwhile, MiniONN [18] designs fast matrix multiplication protocols. Gazelle [6] and Delphi [19] combine techniques from HE and MPC to achieve fast private inference. CrypTFlow2 [24] designs an efficient comparison protocol, and its followup works [41, 42, 40] focus on improving performance. In addition, there exists many works [7, 44, 43] for large language model inference combines techniques from HE and secret sharing. In order to solve the performance bottleneck of 2PC, Chameleon [22] employed a semi-honest third-party to generate correlated randomness for multiplication triplets in offline. Further, for improving performance in non-linear functions such as ReLU, SecureNN [12] constructs novel protocols for comparsion with the help of a third-party. For providing better overall efficiency, some works ABY3 [9], Falcon[27], Swift [30] are propoed based on 3PC replicated secret sharing. Additionally,  [45, 28] and [34, 31, 33, 32] focus on resisting malicious adversaries and require much more costs in 3PC and 4PC, respectively.
    Instead, a critical limitation common to all these representative works lies in their restriction to a maximum of four participating parties, consequently exhibiting inadequate scalability. In order to solve the bottleneck, many works are proposed. Motion [47] designs an nnPC framework that combines secure computation schemes based on arithmetic sharing, boolean sharing, and GC. Howerver, since computation is required between every pair of parties in Motion, the overall efficiency is limited. Alessandro et al. [46] proposes a nnPC framework by utilizing replicated secret sharing, but its sharing structure is relatively complex and incurs high communication overhead for multiplication operations. Liu et al. [5] first constructs a practical inference framework using Shamir secret sharing. Unfortunately, when dealing with deep neural networks such as VGG16, the communication overhead remains substantial. Franklin et al. [1] introduced the PSS and its follow-up works [8, 3, 48] focus on improving multiplication operation and resisting malicious adversaries in theory. In this work, we mainly achieve efficient vector-matrix multiplication, convolution computation and non-linear operations for more practical secure inference using PSS.
@@ VIII Conclusion & Future Work
    In this paper, we construct secure blocks based on packed Shamir secret sharing for high-throughput and scalable secure inference. Extensive evaluations also present our improvements. For future work, we are willing to extend to malicious adversary scenarios and improve computational efficiency by GPU.
    -A Functionality Description
    Functionality $\mathcal{F}_{Pmult-DN}$ Input: The functionality receives inputs $\llbracket\boldsymbol{a}\rrbracket_{d}$ and $\llbracket\boldsymbol{b}\rrbracket_{d}$. Output: Compute the following 1. Reconstruct $\boldsymbol{a}\in\mathbb{F}_{p}^{k}$ and $\boldsymbol{b}\in\mathbb{F}_{p}^{k}$ to compute $\boldsymbol{c}_{i}=\boldsymbol{a}_{i}\boldsymbol{b}_{i},i\in[k]$. 2. Distribute shares of $\llbracket\boldsymbol{c}\rrbracket_{d}$ to the parties.
    Functionality $\mathcal{F}_{PMatMult}^{Trunc}$ Input: The functionality receives inputs $\llbracket A\rrbracket_{d}\in\mathbb{F}_{p}^{u\times v}$ and $\llbracket B\rrbracket_{d}\in\mathbb{F}_{p}^{v\times w}$. Output: Compute the following 1. Reconstruct $A^{i}\in\mathbb{F}_{p}^{u\times v}$ and $B^{i}\in\mathbb{F}_{p}^{v\times w},i\in[k]$ to compute $C^{\prime i}=A^{i}\cdot B^{i},i\in[k]$. 2. Compute $C^{i}=C^{\prime i}_{\alpha,\beta}/2^{\ell_{x}},i\in[k],\alpha\in[u],\beta\in[w]$. 3. Distribute shares of $\llbracket C\rrbracket_{d}\in\mathbb{F}_{p}^{u\times w}$ to the parties.
    Functionality $\mathcal{F}_{VecMatMult}^{Trunc}$ Input: The functionality receives inputs a vector $\llbracket\boldsymbol{a}\rrbracket_{d}\in\mathbb{F}_{q}^{u}$ and a matrix $\llbracket\boldsymbol{A}\rrbracket_{d}\in\mathbb{F}_{q}^{u\times vk}$. Output: Compute the following 1. Reconstruct $\boldsymbol{a}\in\mathbb{F}_{p}^{ku}$ and $\boldsymbol{A}\in\mathbb{F}_{p}^{ku\times vk}.$ 2. Compute $\boldsymbol{c}=\boldsymbol{a}\cdot A$ and compute $\boldsymbol{c}_{i}^{\prime}=\boldsymbol{c}_{i}/2^{\ell_{x}}$. 3. Distribute shares of $\llbracket\boldsymbol{c}^{\prime}\rrbracket_{d}\in\mathbb{F}_{q}^{v}$ to the parties.
    Functionality $\mathcal{F}_{RandomPairs}$ Input: The functionality receives no inputs. Output: Compute the following 1. Randomly sample a vecotr $\boldsymbol{r}\in\mathbb{F}_{p}^{kv}$ and compute a vector $\boldsymbol{r}^{\prime}=(\sum_{i=0}^{k-1}\boldsymbol{r}_{i},\cdots,\sum_{i=(v-1)k}^{kv-1}\boldsymbol{r}_{i})\in\mathbb{F}_{p}^{v}$. 2. Distribute shares of the random pairs $(\llbracket\boldsymbol{r}\rrbracket_{2d},\llbracket\boldsymbol{r^{\prime}}\rrbracket_{d})$ to the parties.
[…]
@@
    Batch generation of random shares using Vandermonde matrices. The protocol implementation in Algorithm 13. It is a natural extension of protocols for generating random shares based on Shamir secret sharing. Note that the degree of the polynomial chosen in Algorithm 13 is $d$. We can select any degree $d_{1}$ as long as it satisfies $d\leq d_{1}\leq n-1$. The amortized communication complexity per generated share is 1 round with $\frac{n-1}{k(n-t)}$ field elements.
@@
    [3] designed a sharing transformation protocol to securely perform the degree transformation. Specifically, given a PSS $\llbracket\boldsymbol{x}\rrbracket_{n-1}$, it outputs $\llbracket\boldsymbol{x}\rrbracket_{n-k}.$ We extend this protocol simply to support arbitrary degree conversions for Shamir secret sharing and PSS in Algorithm 14. Note that Step 1 can be performed during the offline phase. The amortized communication complexity of the protocol is 2 round with $\frac{2(n-1)}{k}(\frac{1}{n-t}+\frac{1}{n})$ field elements (resp. $2(n-1)(\frac{1}{n-t}+\frac{1}{n})$ field elements for Shamir secret sharing) per party.
@@
    [35] proposed a protocol to securely generate the shares of random bits. The parties generate a bits $r\in\{0,1\}$ by computing $S(a)=\frac{\frac{a}{\sqrt{a^{2}}}+1}{2}$, where $a\in\mathbb{F}_{p}^{*}$ is a uniformly random non-zero element. $\sqrt{a^{2}}$ is formulated to equal the unique element in [1,(p1)/2][1,(p-1)/2] and $a/\sqrt{a^{2}}$ is $1$ or 1-1. We extend this protocol using PSS in Algorithm 15. The amortized communication complexity per generated random bits is 4 rounds with $\frac{n-1}{k}(\frac{5}{n-t}+\frac{4}{n}+k+1)$ field elements.
@@
      Compute $\llbracket\boldsymbol{a}^{\prime}\rrbracket_{d}\leftarrow\mathcal{F}_{Pmult-DN}(\llbracket\boldsymbol{a}\rrbracket_{d},\llbracket\boldsymbol{a}\rrbracket_{d})$.
@@
    Given two inputs $\llbracket\boldsymbol{a}\rrbracket_{d_{1}},\llbracket\boldsymbol{b}\rrbracket_{d},k-1\leq d_{1}\leq d$, we get $\llbracket\boldsymbol{a}\rrbracket_{d_{1}}\oplus\llbracket\boldsymbol{b}\rrbracket_{d}$ by computing $\llbracket\boldsymbol{a}\rrbracket_{d_{1}}\oplus\llbracket\boldsymbol{b}\rrbracket_{d}=\llbracket\boldsymbol{a}\rrbracket_{d_{1}}+\llbracket\boldsymbol{b}\rrbracket_{d}-2\llbracket\boldsymbol{a}\rrbracket_{d_{1}}\llbracket\boldsymbol{b}\rrbracket_{d}$, where $\boldsymbol{a},\boldsymbol{b}$ are bit strings. The protocol implementation in Algorithm 16. The amortized communication complexity per bit xor is the same as protocol $\Pi_{DegreeTrans}$.
@@
    -C Proofs
    Theorem 13 (Theorem 4).
@@
    For security proof, we only describe the simulation for protocol $\Pi_{VecMatMult}^{Trunc}$. The security proof for other protocols are simple compositions of local computations and invocations of ideal functionalities.
@@
    Theorem 14 (Theorem 3).
    $\Pi_{VecMatMult}^{Trunc}$ securely realizes $\mathcal{F}_{VecMatMult}^{Trunc}$ in the $\mathcal{F}_{TruncTriple}$-hybrid model with abort, in the presence of a fully semi-honest adversary controlling $t$ corrupted parties.

### v2

@@ VI-C Benchmarks on Secure Inference
    VI-C1 Communication Costs of Secure Inference
@@
    The communication costs of the protocol from LXY24 and our protocol on various datasets and neural networks are shown in Table III. Compared to LXY24, we achieve a significant advantage in terms of offline communication, online communication, and total communication costs. Moreover, our analysis reveals a progressively more pronounced advantage along the diagonal from the upper-left to the lower-right corner of the table. For instance, compared to LXY24, our approach improves the communication $1.3-5.9\times$, $1.6-11.2\times$, and $1.4-6.8\times$ in offline, online, and total communication overhead respectively as $k$ increases for AlexNet. Additionally, for VGG16, LXY24 suffers from memory overflow under both the 31PC where 12 parties are controlled by the adversary and 63PC, whereas our approach continues to execute successfully. This advantage primarily stems from the fact that more parties enable the selection of larger values for $k$ in PSS, where an increased $k$ confers greater benefits for deeper neural networks and larger datasets.
    VI-C2 Running Time of Secure Inference (WAN)
@@
    The running time of LXY24 and ours in WAN is shown in Table V. Similar to Table III, our advantages become increasingly significant from the top-left to the bottom-right of the table. However, it is worth noting that in the case of 11PC, there is no improvement compared to LXY24. This phenomenon primarily stems from our scheme is significantly affected by computational overhead, and when a small $k$ is selected, the communication efficiency improvement is limited. Furthermore, we can find that when the communication overhead increases by more than $3.8\times$, our online running time will be significantly improved. For example, the online phase is $2.4-2.6\times$ faster under 63PC for AlexNet. At the same time, as the number of parties and $k$ increase, the offline running time and the total running time can be improved by $1.1-1.6\times$ and $1.1-1.8\times$ respectively. Additionally, ours performs significantly better for deeper networks. For instance, the performance on AlexNet is better than that on MiniONN. The reason is that deeper networks involve more convolution operations, and the improvements in our communication costs and running time become more pronounced. In section VI-B, it can be observed that our method achieves significant advantages in convolution operations.
    VI-C3 Running Time of Secure Inference (LAN)
@@
    The running time of LXY24 and ours in LAN is shown in Table IV. We observe that in the online phase, our running time shows little to no advantage compared to LXY24, as our scheme is more heavily affected by computational overhead. However, the offline and the total running time, as the number of parties and $k$ increases, will also have significant improvements, with the increase being $1.1-3.8\times$ and $1.2-2.6\times$ respectively.
@@
    In a word, our improvements in the LAN setting are less pronounced compared to those in the WAN setting. The reason lies in the higher computational overhead required by our scheme. Fortunately, as the number of parties and $k$ increases, our scheme exhibits more substantial advantages, mainly owing to our efficient vector-matrix multiplication, parallel convolution computation, and parallel non-linear computation based on PSS. Moreover, MPC protocols are more commonly deployed in WAN settings, so our improvements hold practical significance.
@@ VII Related Work
    Privacy-preserving deep learning has been a prominent research focus in recent years. Based on the number of participating parties, existing works can be broadly categorized into two categories: those involving a small number of parties (e.g., 2PC, 3PC, 4PC) and those designed for a large number of parties (nnPC).
    In the area of two-party computation (2PC), ABY [6] and ABY2.0 [7] design a framework that combines secure computation schemes based on arithmetic sharing , boolean sharing, and Yao’s garbled circuits (GC). SecureML [5] proposes a scheme in the two-server outsourced computation models with secret sharing and GC. Meanwhile, MiniONN [9] designs fast matrix multiplication protocols. Gazelle [8] and Delphi [36] combine techniques from HE and MPC to achieve fast private inference. CrypTFlow2 [12] designs an efficient comparison protocol, and its followup works [37, 38, 39] focus on improving performance. In addition, there exists many works [40, 41, 42] for large language model inference combines techniques from HE and secret sharing. In order to solve the performance bottleneck of 2PC, Chameleon [10] employed a semi-honest third-party to generate correlated randomness for multiplication triplets in offline. Further, for improving performance in non-linear functions such as ReLU, SecureNN [16] constructs novel protocols for comparsion with the help of a third-party. For providing better overall efficiency, some works ABY3 [15], Falcon[17], Swift [20] are propoed based on 3PC replicated secret sharing. Additionally,  [43, 18] and [24, 21, 23, 22] focus on resisting malicious adversaries and require much more costs in 3PC and 4PC, respectively.
    Instead, a critical limitation common to all these representative works lies in their restriction to a maximum of four participating parties, consequently exhibiting inadequate scalability. In order to solve the bottleneck, many works are proposed. Motion [44] designs an nnPC framework that combines secure computation schemes based on arithmetic sharing, boolean sharing, and GC. Howerver, since computation is required between every pair of parties in Motion, the overall efficiency is limited. Alessandro et al. [45] propose an nnPC framework by utilizing replicated secret sharing, but its sharing structure is relatively complex and incurs high communication overhead for multiplication operations. LXY24 first constructs a practical inference framework using Shamir secret sharing. Unfortunately, when dealing with deep neural networks such as VGG16, the communication overhead remains substantial. Franklin et al. [27] introduced the PSS and its follow-up works [46, 29, 47] focus on improving multiplication operation and resisting malicious adversaries in theory. In this work, we mainly achieve efficient vector-matrix multiplication, convolution computation and non-linear operations for more practical PPDL using PSS.
@@ VIII Conclusion & Future Work
    In this paper, we construct secure blocks based on packed Shamir secret sharing for high-throughput and scalable privacy-preserving deep learning. Extensive evaluations also present our improvements. For future work, we are willing to extend to malicious adversary scenarios and improve computational efficiency by GPU.
    -A The Offline Costs of Microbenchmarks
    The offline costs of the main sub-protocols are given in Table VI. Concretely, for linear operations (i.e., convolution and matrix multiplication operations), there is no advantage in running time. The reason is that the offline complexity for generating random bits and truncation triples is greater. For example, generating truncation triples needs to call the degree transformation protocol (Algorithm 14) multiple times. However, it is worth noting that our communication costs still show improvement with a larger $k$. Take the case with $k=8$ as an example, the offline cost of matrix multiplication achieves improvement. Therefore, we can reasonably infer that when a larger $k$ is chosen, the communication cost is further reduced, which in turn leads to an advantage in running time. This is reflected in the Table V. For the non-linear operations (i.e., ReLU and Maxpool), we reduce the communication costs by $3.1-3.5\times$ compared to LXY24. Furthermore, we are $1.5-2.0\times$ faster in WAN and $1.1-1.6\times$ faster in LAN.
    TABLE VI: The offline running time and amortized communication costs of per party of the sub-protocol. Communication is in MB. Running time is in seconds. $n=21$ is the number of parties. $t$ is the number of adversaries. $k$ is the size of the package in PSS. $\mathrm{Conv}_{32,32,64,5,1}$$\mathrm{MatMult}_{16,128,8192}$$\mathrm{ReLU}_{64}$$\mathrm{Maxpool}_{16,8,2}$.
    Operation t Protocol k Comm.(MB) WAN(s) LAN(s)
    $\mathrm{Conv}$ 3 LXY24 - 65.74 65.66 7.54
    Ours 8 62.12 72.28 10.67
    7 LXY24 - 77.18 72.41 10.04
    Ours 4 76.29 83.19 11.92
    $\mathrm{MatMult}$ 3 LXY24 - 141.94 140.27 16.32
    Ours 8 117.49 203.83 32.37
    7 LXY24 - 165.97 154.22 21.49
    Ours 4 205.77 225.00 33.02
    $\mathrm{ReLU}$ 3 LXY24 - 26.81 15.73 2.47
    Ours 8 7.68 8.49 1.56
    7 LXY24 - 23.49 16.50 2.44
    Ours 4 11.04 11.24 2.17
    $\mathrm{Maxpool}$ 3 LXY24 - 35.34 22.31 2.92
    Ours 8 10.12 11.12 2.01
    7 LXY24 - 30.97 21.48 3.74
    Ours 4 14.55 14.82 2.53
@@
    Batch generation of random shares using Vandermonde matrices. The protocol implementation in Algorithm 13. It is a natural extension of protocols for generating random shares based on Shamir secret sharing. Note that the degree of the polynomial chosen in Algorithm 13 is $d$. We can select any degree $d_{1}$ as long as it satisfies $d\leq d_{1}\leq n-1$. The communication complexity is 1 round with $\frac{n-1}{k(n-t)}$ field elements amortized per generated share.
@@
    [29] designed a sharing transformation protocol to securely perform the degree transformation. Specifically, given a PSS $\llbracket\boldsymbol{x}\rrbracket_{n-1}$, it outputs $\llbracket\boldsymbol{x}\rrbracket_{n-k}.$ We extend this protocol simply to support arbitrary degree conversions for Shamir secret sharing and PSS in Algorithm 14. Note that Step 1 can be performed during the offline phase. The communication complexity of the protocol is 2 round with $\frac{2(n-1)}{k}(\frac{1}{n-t}+\frac{1}{n})$ field elements amortized per party (resp. $2(n-1)(\frac{1}{n-t}+\frac{1}{n})$ field elements for Shamir secret sharing).
@@
    [48] proposed a protocol to securely generate the shares of random bits. The parties generate a bits $r\in\{0,1\}$ by computing $S(a)=\frac{\frac{a}{\sqrt{a^{2}}}+1}{2}$, where $a\in\mathbb{F}_{p}^{*}$ is a uniformly random non-zero element. $\sqrt{a^{2}}$ is formulated to equal the unique element in [1,(p1)/2][1,(p-1)/2] and $a/\sqrt{a^{2}}$ is $1$ or 1-1. We extend this protocol using PSS in Algorithm 15. The communication complexity is 4 rounds with $\frac{n-1}{k}(\frac{5}{n-t}+\frac{4}{n}+k+1)$ field elements amortized per generated random bits.
@@
      Compute $\llbracket\boldsymbol{a}^{\prime}\rrbracket_{d}\leftarrow\mathcal{F}_{PDN}(\llbracket\boldsymbol{a}\rrbracket_{d},\llbracket\boldsymbol{a}\rrbracket_{d})$.
@@
    Given two inputs $\llbracket\boldsymbol{a}\rrbracket_{d_{1}},\llbracket\boldsymbol{b}\rrbracket_{d},k-1\leq d_{1}\leq d$, we get $\llbracket\boldsymbol{a}\rrbracket_{d_{1}}\oplus\llbracket\boldsymbol{b}\rrbracket_{d}$ by computing $\llbracket\boldsymbol{a}\rrbracket_{d_{1}}\oplus\llbracket\boldsymbol{b}\rrbracket_{d}=\llbracket\boldsymbol{a}\rrbracket_{d_{1}}+\llbracket\boldsymbol{b}\rrbracket_{d}-2\llbracket\boldsymbol{a}\rrbracket_{d_{1}}\llbracket\boldsymbol{b}\rrbracket_{d}$, where $\boldsymbol{a},\boldsymbol{b}$ are bit strings. The protocol implementation in Algorithm 16. The communication complexity is the same as protocol $\Pi_{DegreeTrans}$ amortized per xor operation.
@@
    -C Functionality Description
    Functionality $\mathcal{F}_{PDN}$ Input: The functionality receives inputs $\llbracket\boldsymbol{a}\rrbracket_{d}$ and $\llbracket\boldsymbol{b}\rrbracket_{d}$. Output: Compute the following 1. Reconstruct $\boldsymbol{a}\in\mathbb{F}_{p}^{k}$ and $\boldsymbol{b}\in\mathbb{F}_{p}^{k}$ to compute $\boldsymbol{c}_{i}=\boldsymbol{a}_{i}\boldsymbol{b}_{i},i\in[k]$. 2. Distribute shares of $\llbracket\boldsymbol{c}\rrbracket_{d}$ to the parties.
    Functionality $\mathcal{F}_{PMatMult}^{Fixed}$ Input: The functionality receives inputs $\llbracket A\rrbracket_{d}\in\mathbb{F}_{p}^{u\times v}$ and $\llbracket B\rrbracket_{d}\in\mathbb{F}_{p}^{v\times w}$. Output: Compute the following 1. Reconstruct $A^{i}\in\mathbb{F}_{p}^{u\times v}$ and $B^{i}\in\mathbb{F}_{p}^{v\times w},i\in[k]$ to compute $C^{\prime i}=A^{i}\cdot B^{i},i\in[k]$. 2. Compute $C^{i}=C^{\prime i}_{\alpha,\beta}/2^{\ell_{x}},i\in[k],\alpha\in[u],\beta\in[w]$. 3. Distribute shares of $\llbracket C\rrbracket_{d}\in\mathbb{F}_{p}^{u\times w}$ to the parties.
    Functionality $\mathcal{F}_{VecMatMult}^{Fixed}$ Input: The functionality receives inputs a vector $\llbracket\boldsymbol{a}\rrbracket_{d}\in\mathbb{F}_{q}^{u}$ and a matrix $\llbracket\boldsymbol{A}\rrbracket_{d}\in\mathbb{F}_{q}^{u\times vk}$. Output: Compute the following 1. Reconstruct $\boldsymbol{a}\in\mathbb{F}_{p}^{ku}$ and $\boldsymbol{A}\in\mathbb{F}_{p}^{ku\times vk}.$ 2. Compute $\boldsymbol{c}=\boldsymbol{a}\cdot A$ and compute $\boldsymbol{c}_{i}^{\prime}=\boldsymbol{c}_{i}/2^{\ell_{x}}$. 3. Distribute shares of $\llbracket\boldsymbol{c}^{\prime}\rrbracket_{d}\in\mathbb{F}_{q}^{v}$ to the parties.
    Functionality $\mathcal{F}_{VM-RandTuple}$ Input: The functionality receives no inputs. Output: Compute the following 1. Randomly sample a vecotr $\boldsymbol{r}\in\mathbb{F}_{p}^{kv}$ and compute a vector $\boldsymbol{r}^{\prime}=(\sum_{i=0}^{k-1}\boldsymbol{r}_{i},\cdots,\sum_{i=(v-1)k}^{kv-1}\boldsymbol{r}_{i})\in\mathbb{F}_{p}^{v}$. 2. Distribute shares of the random pairs $(\llbracket\boldsymbol{r}\rrbracket_{2d},\llbracket\boldsymbol{r^{\prime}}\rrbracket_{d})$ to the parties.
[…]
    -D Proofs
    Proposition 1.
    If $xy\leq 2^{k}$ for some $k<\ell$, then with probability at least $1-2^{k}/(2^{\ell}-1)$ it holds that $z=\lfloor xy/2^{\ell_{x}}\rfloor+v$ for some $v\in\{0,1\}$.
    Proof.
    Let $z=xy,z^{\prime}=(z+r)/2^{\ell_{x}}-r^{\prime},$
    $z^{\prime}$ $=\frac{z+r-((z+r)\pmod{2^{\ell_{x}}})}{2^{\ell_{x}}}-r^{\prime}$ (1)
    $=\frac{z+r-(2^{\ell}-1)u-(z\pmod{2^{\ell_{x}}}+r\pmod{2^{\ell_{x}}})}{2^{\ell_{x}}}$ (2)
    $+\frac{2^{\ell_{x}}v}{2^{\ell_{x}}}-r^{\prime}$ (3)
    $=\frac{z-z\pmod{2^{\ell_{x}}}}{2^{\ell_{x}}}+\frac{r-r\pmod{2^{\ell_{x}}}}{2^{\ell_{x}}}$ (4)
    $-r^{\prime}+v-\frac{(2^{\ell}-1)u}{2^{\ell_{x}}}$ (5)
    $=z/2^{\ell_{x}}+r^{\prime}+v-r^{\prime}-\frac{(2^{\ell}-1)u}{2^{\ell_{x}}}$ (6)
    $=z/2^{\ell_{x}}+v-2^{{\ell}-{\ell_{x}}}u,$ (7)
    where $u=((z+r)\geq p),p=2^{\ell}-1,v=((z^{\prime}\pmod{2^{d}}+r^{\prime}\pmod{2^{\ell_{x}}})\geq 2^{\ell_{x}}).$ The error is $v-2^{{\ell}-{\ell_{x}}}u$. The probability that $u=1$ is the probability that $r\geq p-z$. Moreover, since $z<2^{k}$ and $r$ is uniformly random in $\mathbb{F}_{p}$, this probability is upper bounded by $2^{k}/(2^{\ell}-1)$. ∎
    Proposition 2.
@@
    For security proof, we only describe the simulation for protocol $\Pi_{VecMatMult}^{Fixed}$. The security proof for other protocols are simple compositions of local computations and invocations of ideal functionalities.
@@
    Theorem 1.
    $\Pi_{VecMatMult}^{Fixed}$ securely realizes $\mathcal{F}_{VecMatMult}^{Fixed}$ in the $\mathcal{F}_{TruncTriple}$-hybrid model with abort, in the presence of a fully semi-honest adversary controlling $t$ corrupted parties.